OpenStudy (anonymous):

calculate the wavelength in nm of a photon emitted by a hydrogen atom when its electron drops from the n=5 state to the n=2 state

OpenStudy (dan815):

use the rydberg eqn

OpenStudy (abb0t):

Use the \(\sf \color{green}{Rhydberg~Equation}\), which is: \(\sf \large\color{red}{\frac{1}{\lambda} = RZ^2(\frac{1}{n_1 ^2}-\frac{1}{n_2 ^2})}\) where \(\sf \lambda\) = wavelegnth of photon R = rhydberg constant \(\sf \color{violet}{1.098 \times 10^7 m^{-1}}\) Z = atomic number of atom \(\sf n_1~and~n_2\) = integers, where \(\sf n_2 > n_1\)